# 2D implementation of Kinetic-diffusion Monte Carlo in Eiron

### v1

@@ 1 Introduction
A variety of hybrid approaches exist that combine kinetic simulation with diffusive simulation [3], including approaches that combine Monte Carlo simulation with deterministic methods, e.g. [2, 8]. However, we make use of asymptotic-preserving Monte Carlo schemes (APMC) [15, 5], which work fully at the Monte Carlo level. Specifically, we make use of the KDMC method [15]. We note that these schemes are a special case of a broader field of asymptotic-preserving schemes, see [9] and references therein.
The general idea behind APMC schemes is that they simulate particle trajectories through time-stepping that alternates between kinetic and diffusive positional increments, weighted in such a way that they produce consistent simulations in both low-collisional (kinetic) regimes and high-collisional (diffusive) regimes. We note that these schemes, in general, do not provide any error guarantees in the intermediate regimes. Any errors in these regimes can however be removed by the combination of APMC with multilevel Monte Carlo techniques [6, 7], see [17, 12, 11]. In the cited works, the APMC approach is shown to significantly reduce the required computational effort for simulating simplified mathematical test problems.
@@ 2.2 Kinetic-diffusion Monte Carlo
We refer to [15] for further details and a full analysis of the scheme, but make the following observations in terms of the charge-exchange rate $R_{\text{cx}}$ in relation to the KDMC discretization time step size $\Delta t$. In low-collisional regimes, i.e., $R_{\text{cx}}\ll\Delta t$, observe that, in general, $\tau\gg\theta$. Hence, the kinetic component (line 4) of Algorithm 2 will dominate and the results produced by Algorithm 2 will converge to those produced by Algorithm 1. Inversely, in high-collisional regimes, i.e., $R_{\text{cx}}\gg\Delta t$, we have that, in general, $\tau\ll\theta$. Hence, the diffusive component (line 8) of Algorithm 2 will dominate. An intuitive explanation for the introduction of the diffusive step in line 8 is that one substitutes the sum of an infinite number of i.i.d. steps of the form given in line 4 by a single normally distributed increment, making use of the law of large numbers.
@@ 4 Simulation results
Throughout this Section, our test-case is a two-dimensional slab with ionizing boundaries. In what follows, we normalize all simulation parameters, assuming that the domain has dimensions $1\;\text{m}\times 1\;\text{m}$. We initialize all particles using an isotropic point-source at the center of the domain with velocities sampled from a 2D-Maxwellian. We then simulate the particles to a given end time $t_{\text{end}}$, at which point we construct a histogram of their final location in the domain using a $128\times 128$-grid. We consider homogeneous collision backgrounds, with a fixed charge-exchange rate $R_{\text{cx}}$. Post-collisional velocities are sampled from an unbiased isotropic Maxwellian. We elaborate on precise numerical parameter values when discussing individual test-cases.
@@ 4.1 Kinetic regime
To visualize the results, we show a KDMC simulation for $\Delta t=2^{-4}\;\text{s}$ in Figure 1, together with a cell-by-cell difference with the reference kinetic simulation. As the two simulations are visually indistinguishable, we opt not to show the reference separately. Comparing the computed solution with the difference, we see that we observe that we achieve between three and four digits of accuracy in the region of interest at the center of the domain.
In Figure 2, we consider the convergence of the KDMC simulation. To reduce variance, we flatten the solution to one dimension, by averaging over the x-dimension. We then mirror the solution around the source at the middle of the domain. In Figure 2(a), we plot the 2-norm of the differences of these curves. We observe that this two norm decreases with roughly an order 3 until it hits a plateau caused by the sampling error due to the finite number of trajectories. This indicates a convergence that is roughly one and a half orders better than that predicted by the theoretical results in [15]. Next to the fact that there is no direct link between the Wasserstein metric on the particles’ location and the 2-norm on the histogram, it is likely that the symmetry of the problem is a cause for this increased rate of convergence. In Figure 2(b), we show the corresponding run times and observe that the KDMC and kinetic simulations incur a comparable computational cost, with KDMC slightly outperforming the kinetic simulation.
@@ 4.2 Diffusive regime
When increasing the collision rate, the modeling error incurred by approximating kinetic increments by diffusion in the KDMC scheme should vanish. In [15], it was show that this convergence should occur as $\mathcal{O}(R_{\text{cx}}^{-1.5})$, in terms of Wasserstein-1 distance of the particle distribution, as $R_{\text{cx}}\to\infty$. To verify the limit we perform simulations with a similar particle source with mean speed $0.0625\;\text{ms}^{-1}$. We fix $\Delta t=1\;\text{s}$ and consider combinations the mean post-collisional speed $\frac{1}{\varepsilon}\sqrt{\frac{\pi}{10}}\frac{1}{512}\;\text{ms}^{-1}$ and collision rate $R_{\text{cx}}=\frac{1}{\varepsilon^{2}}\frac{1}{128}\;\text{s}^{-1}$ with $\varepsilon$ taking values in the range $2^{-7.5},1]$. In this case we use $2\cdot 10^{9}$ particles until time $t_{\text{end}}=4\;\text{s}$. We then generate the similar figures to those in Section  4.1.
@@
(a) Pointwise difference for a selection of values of $\varepsilon$
@@
$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$R_{\text{cx}}$ $\text{s}^{-1}$]Runtime [s]KineticKDMC
@@
In Figure 3, we show the KDMC simulation and pointwise error for the case $R_{\text{cx}}=256\;\text{s}^{-1}$ and the mean post-collisional speed $0.19817\;\text{ms}^{-1}$. Note that the shape of the solution differs from that for the kinetic case, in that it resembles a bell-curve, as opposed to the exponential decay from the center source, observed in Figure 1. We consider the convergence behavior in Figure 4. On the right-hand side of Figure 4(b), we empirically observe a convergence rate of with approximate order 1 as $R_{\text{cx}}$ increases. Notably, this convergence is slower than the known bound in the Wasserstein-1 distance. However, in parallel work [21] a similar rate was shown in the $L_{2}$-norm, when computing time-integrated quantities using KDMC, combined with the diffusive estimator presented in [16].
On the right-hand side of the figure, we see a similar phenomenon to that shown in Figure 2(a), i.e., as $R_{\text{cx}}$ decreases, the both simulations converge to kinetic simulations until a plateau is achieved due to the sampling error. What is initially surprising, is that we observe a sharp dip approximately $R_{\text{cx}}=4\;\text{s}^{-1}$. To explain this phenomenon, we plot the pointwise differences of the averaged and folded simulation results for a selection of values of $R_{\text{cx}}$ in Figure 4(a). In this figure, we see that KDMC underestimates the particles’ displacement in kinetic regimes and overestimates the displacement in diffusive regimes (compare the curves for $R_{\text{cx}}=2\;\text{s}^{-1}$ [peak error on the kinetic side of Figure 4(b)] and $R_{\text{cx}}=11.314\;\text{s}^{-1}$ [peak error on the diffusive side of Figure 4(b)]). Hence, during the transition between the two regimes, there is a value of $R_{\text{cx}}$ for which these two errors cancel, causing the sharp dip in error.
In Figure 4(c), we show that KDMC achieves a significant speedup in the diffusive regime, growing proportionally to $R_{\text{cx}}$. Here it is clear that the runtime of the kinetic simulations directly scale with the number of collision events that need to be resolved, while that for the KDMC simulations stays more-or-less constant. In the highest-collisional cases, this results in a speedup of between two and three orders of magnitude, showing the strength of KDMC over standard kinetic simulation in this regime.
@@ 5 Conclusions
We have introduced the first two-dimensional implementation of the KDMC scheme for the Boltzman-BGK equation in the Eiron code. Using this implementation, we confirmed that the results produced by KDMC converge to those produced by a standard kinetic simulation in both the kinetic limit ($\Delta t\to 0$) and diffusive limit ($R_{\text{cx}}\to\infty$). Our measured 2-norm convergence, does not match the rates given in [15] for the Wasserstein-1 distance. However, we believe the improved rate in $\Delta t$ to be due to the symmetry in the considered test-case and our rate in the limit $R_{\text{cx}}\to 0$ matches the $L_{2}$-norm results shown in parallel work on analyzing time-integrated simulations with KDMC. In the kinetic regime, KDMC simulations attain similar run times to the equivalent kinetic simulations. However, in diffusive regimes, they outperform the kinetic reference simulations. In some cases, by multiple orders of magnitude.
@@ Acknowledgments
The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources. Emil Løvbak was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 563450842. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.

### v3

@@ 1 Introduction
A variety of hybrid approaches exist that combine kinetic simulation with diffusive simulation [3], including approaches that combine Monte Carlo simulation with deterministic methods, e.g. [2, 8]. However, we make use of asymptotic-preserving Monte Carlo schemes (APMC) [15, 5], which work fully at the Monte Carlo level. Specifically, we make use of the kinetic-diffusion Monte Carlo method (KDMC) [15]. We note that these schemes are a special case of a broader field of asymptotic-preserving schemes, see [9] and references therein.
The general idea behind APMC schemes is that they simulate particle trajectories through time-stepping that alternates between kinetic and diffusive positional increments in such a way that mainly simulate kinetic dynamics in low-collisional (kinetic) regimes, and diffusive dynamics in high-collisional (diffusive) regimes. We note that these schemes, in general, do not provide any error guarantees in the intermediate regimes. Any errors in these regimes can however be removed by the combination of APMC with multilevel Monte Carlo techniques [6, 7], see [17, 12, 11]. In the cited works, the APMC approach is shown to significantly reduce the required computational effort for simulating simplified mathematical test problems.
@@ 2.2 Kinetic-diffusion Monte Carlo
We refer to [15] for further details and a full analysis of the scheme, but make the following observations in terms of the charge-exchange rate $R_{\text{cx}}$ in relation to the KDMC discretization time step size $\Delta t$. In low-collisional regimes, i.e., $R_{\text{cx}}\ll\Delta t^{-1}$, observe that, in general, $\tau\gg\theta$. Hence, the kinetic component (line 4) of Algorithm 2 will dominate and the results produced by Algorithm 2 will converge to those produced by Algorithm 1. Inversely, in high-collisional regimes, i.e., $R_{\text{cx}}\gg\Delta t^{-1}$, we have that, in general, $\tau\ll\theta$. Hence, the diffusive component (line 8) of Algorithm 2 will dominate. An intuitive explanation for the introduction of the diffusive step in line 8 is that one substitutes the sum of an infinite number of i.i.d. steps of the form given in line 4 by a single normally distributed increment, making use of the law of large numbers.
@@ 4 Simulation results
Throughout this Section, our test-case is a two-dimensional slab with ionizing boundaries. In what follows, we normalize all simulation parameters, assuming that the domain has dimensions $1\;\text{m}\times 1\;\text{m}$. We initialize all particles using an isotropic point-source at the center of the domain with velocities sampled from a 2D-Maxwellian. We then simulate the particles to a given end time $t_{\text{end}}$, at which point we construct a scaled histogram of their final location in the domain using a $128\times 128$-grid, such that the solution integrates to one. We consider homogeneous collision backgrounds, with a fixed charge-exchange rate $R_{\text{cx}}$. Post-collisional velocities are sampled from an unbiased isotropic Maxwellian. We elaborate on precise numerical parameter values when discussing individual test-cases.
@@ 4.1 Kinetic regime
To visualize the results, we show a KDMC simulation for $\Delta t=2^{-4}\;\text{s}$ in Figure 1, together with a cell-by-cell difference with the reference kinetic simulation. As the two simulations are visually indistinguishable, we opt not to show the reference separately. Comparing the computed solution with the difference, we observe that we achieve between three and four digits of accuracy in the region of interest at the center of the domain.
In Figure 2, we consider the convergence of the KDMC simulation. To reduce variance, we flatten the solution to one dimension, by averaging over the x-dimension. We then mirror the solution around the source at the middle of the domain. In Figure 2(a), we plot the 2-norm of the differences of these curves. We observe that this two norm decreases with roughly an order 3 until it hits a plateau caused by the sampling error due to the finite number of trajectories. This indicates a convergence that is roughly one and a half orders better than that predicted by the theoretical results in [15]. Next to the fact that there is no direct link between the Wasserstein metric on the particles’ location and the 2-norm on the histogram, it is likely that the symmetry of the problem is a cause for this increased rate of convergence. In Figure 2(b), we show the corresponding run times and observe that the KDMC and kinetic simulations incur a comparable computational cost in the kinetic regime.
@@ 4.2 Diffusive regime
When increasing the collision rate, the modeling error incurred by approximating kinetic increments by diffusion in the KDMC scheme should vanish. In [15], it was show that this convergence should occur as $\mathcal{O}(R_{\text{cx}}^{-1.5})$, in terms of Wasserstein-1 distance of the particle distribution, as $R_{\text{cx}}\to\infty$. To verify the limit we perform simulations with a similar particle source with mean speed $0.0625\;\text{ms}^{-1}$. We fix $\Delta t=1\;\text{s}$ and consider combinations the mean post-collisional speed $\frac{1}{\varepsilon}\sqrt{\frac{\pi}{10}}\frac{1}{512}\;\text{ms}^{-1}$ and collision rate $R_{\text{cx}}=\frac{1}{\varepsilon^{2}}\frac{1}{128}\;\text{s}^{-1}$ with $\varepsilon$ taking values in the range $2^{-7.5},1]$. In this case we use $2\cdot 10^{9}$ particles until time $t_{\text{end}}=4\;\text{s}$. We then generate similar figures to those in Section  4.1.
@@
(a) Pointwise difference for a selection of values of $R_{\text{cx}}$
@@
$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$10^{2}$$10^{3}$$10^{4}$$R_{\text{cx}}$ $\text{s}^{-1}$]Runtime [s]KineticKDMC
@@
In Figure 3, we show the KDMC simulation and pointwise error for the case $R_{\text{cx}}=256\;\text{s}^{-1}$ and the mean post-collisional speed $0.19817\;\text{ms}^{-1}$. Note that the shape of the solution differs from that for the kinetic case, in that it resembles a bell-curve, as opposed to the exponential decay observed in Figure 1, which was faster closer to the center source but slower further away. We consider the convergence behavior in Figure 4. In Figure 4(b), we empirically observe a convergence rate of with approximate order 1 for $R_{\text{cx}}\geq 16\;\text{s}^{-1}$. Notably, this convergence is slower than the known bound in the Wasserstein-1 distance. However, in parallel work [21] a similar rate was shown in the $L_{2}$-norm, when computing time-integrated quantities using KDMC, combined with the diffusive estimator presented in [16].
For $R_{\text{cx}}\leq 2\;\text{s}^{-1}$, we see a similar phenomenon in Figure 4(b) to that shown in Figure 2(a), i.e., as $R_{\text{cx}}$ decreases, the both simulations converge to kinetic simulations until a plateau is achieved due to the sampling error. In the intermediate regime, the relative 2-norm of the differences achieves a maximum of less than 1%. What is initially surprising, is that we observe a sharp dip approximately $R_{\text{cx}}=4\;\text{s}^{-1}$. To explain this phenomenon, we plot the pointwise differences of the averaged and folded simulation results for a selection of values of $R_{\text{cx}}$ in Figure 4(a). In this figure, we see that the error curves for the averaged and flattened results change sign at one point (compare the curves for $R_{\text{cx}}=2\;\text{s}^{-1}$ [peak error on the kinetic side of Figure 4(b)] and $R_{\text{cx}}=11.314\;\text{s}^{-1}$ [peak error on the diffusive side of Figure 4(b)]). At the point of transition, where the sign of the error changes  (roughly $R_{\text{cx}}=4\;\text{s}^{-1}$), there is therefore an artificially reduced error.
In Figure 4(c), we show that KDMC achieves a significant speedup in the diffusive regime, growing proportionally to $R_{\text{cx}}$. Here it is clear that the runtime of the kinetic simulations directly scale with the number of collision events that need to be resolved, while that for the KDMC simulations stays more-or-less constant. In the highest-collisional cases, this results in a speedup of two orders of magnitude, showing the strength of KDMC over standard kinetic simulation in this regime.
@@ 5 Conclusions
We have introduced the first two-dimensional implementation of the KDMC scheme for the Boltzman-BGK equation in the Eiron code. Using this implementation, we confirmed that the results produced by KDMC converge to those produced by a standard kinetic simulation in both the kinetic limit ($\Delta t\to 0$) and diffusive limit ($R_{\text{cx}}\to\infty$). Our measured 2-norm convergence, does not match the rates given in [15] for the Wasserstein-1 distance. However, we believe the improved rate in $\Delta t$ to be due to the symmetry in the considered test-case and our rate in the limit $R_{\text{cx}}\to 0$ matches the $L_{2}$-norm results shown in parallel work on analyzing time-integrated simulations with KDMC. The maximum observed relative error on the averaged and folded results is less that 1% in the considered test cases. Care should be taken in generalizing this observation beyond these test cases as they do not contain, e.g., time averaging source-term estimation procedures such as those used in practical scrape-off layer simulations. However, we note that multilevel Monte Carlo approaches inspired by [17, 12, 11] can be leveraged for further error reduction in these regions if needed. In the kinetic regime, KDMC simulations attain similar run times to the equivalent kinetic simulations. However, in diffusive regimes, they outperform the kinetic reference simulations. In some cases, by multiple orders of magnitude.
@@ Acknowledgments
We thank the anonymous reviewers for their efforts in improving this manuscript. The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources. Emil Løvbak was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 563450842. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.